OpenStudy (codyhelton75):

Which of the following is an example of the associative property? a 2(5 - 3) = 10 - 6 b -7 + (19 + 5) = (-7 + 19) + 5 c11 + (6 + 8) = 11 + (28 + 6) d 1/3 • 3 = 1

OpenStudy (zzr0ck3r):

Associative property says \(a+b)+c=a+(b+c)\) So in other words, if I need to add three numbers, say 2,3,4 then it does not matter if I do 2+3 and then add the result to 4 or if a add 2 to the result of (3+4)

OpenStudy (sylph):

Okay, so when you associate yourself with someone, you and that other person would make a "pair", correct?

OpenStudy (codyhelton75):

yeah

OpenStudy (sylph):

Numbers can associate themselves with one another in the same way, forming "pairs". For numbers, however the only way they can identify as a pair is if they are both in parentheses.

OpenStudy (codyhelton75):

ok.next part please

OpenStudy (sylph):

During addition, it doesn't matter what order the numbers are in, because they all add up to the same digits. \(6+4+8+1=19,~8+1+4+6=19,~etc.\)

OpenStudy (sylph):

The numbers associated with one another don't affect the sum either. \[9+(4+3)=16,~(3+9)+4=16\]

OpenStudy (sylph):

Which of the choices shows that same concept?

OpenStudy (codyhelton75):

b?*guesses nervously*

OpenStudy (codyhelton75):

so?

OpenStudy (sylph):

Yes :)